### v2

# Proxy Prompt: Endowing SAM & SAM 2 with Auto-Interactive-Prompt for Medical Segmentation

Abstract
In this paper, we aim to address the unmet demand for automated prompting and enhanced human-model interactions of SAM and SAM2 for the sake of promoting their widespread clinical adoption. Specifically, we propose Proxy Prompt (PP), auto-generated by leveraging non-target data with a pre-annotated mask. We devise a novel 3-step context-selection strategy for adaptively selecting the most representative contextual information from non-target data via vision mamba and selective maps, empowering the guiding capability of non-target image-mask pairs for segmentation on target image/video data. To reinforce human-model interactions in PP, we further propose a contextual colorization module via a dual-reverse cross-attention to enhance interactions between target features and contextual-embedding with amplifying distinctive features of user-defined object(s). Via extensive evaluations, our method achieves state-of-the-art performance on five public datasets and yields comparable results with fully-trained models, even when trained with only 16 image masks.
1 Introduction
Under the paradigm shifts in Large-scale Vision Models (LVMs), Segment Anything Model (SAM) [16] and SAM 2 [25] have been introduced as generalized foundation models for segmenting and tracking any objects on image and video data, respectively. These models provide a certain degree of interactive segmentation capacity as users can segment any target object(s) according to their needs by using a single model in-one-go. Such capabilities are achieved by leveraging the concept of prompts [32], such as points, boxes, or masks, waiving the traditional demand for massive manually annotations. Instead, users are only required to input prompts directly on target images or video frames.
Figure 1: Illustration of comparison without/with PP in (a-b) SAM 2 using real-time ultrasound frames of 1 subject; and (c-d) SAM using a fundus retina dataset of 3 subjects.
Notwithstanding, widespread clinical adoptions of these LVMs have been substantially impeded by the soaring medical demands for “automated prompting” and “high-level human-model interactions” when it comes to downstream medical tasks, particularly real-time imaging-guided interventions. The current prompting strategies are sub-optimal for two key reasons. First, medical image/video data entails an overwhelmingly huge variations in terms of complexity of the target object(s) to be segmented; segmenting such structures (e.g. vessels) using existing prompts can be practically challenging. As illustrated in Fig. 1(c), for instance, manually inputting either box prompt (Case 1) or point prompt (Case 3) generates poor results; while adopting mask prompt (Case 2) performs well, it is highly tedious, exhaustive and knowledge-demanding for precise mask prompt formation. Second, users are required to input prompt for every single target image/video frame, which is a manual trial-and-error process, tedious, and not user-friendly. Clinical burden becomes exceedingly prominent when segmenting intricated structures Fig. 1(c) and/or massive datasets, especially in resource-limited clinics. Therefore, there is a pressing demand for automated prompt generation to accommodate various clinical needs.
Figure 2: Schematic differences of traditional prompt encoder and our proposed PPG in SAM 2 (a-b) and SAM (c-d).
Apart from this, the existing SAM-based models [27, 41, 38] are inadequate to support the growing demand of high-level human-model interaction to accommodate multifarious clinical goals and high disparity in preferences of clinical users. These models segment target object(s) via training on specific single/multiple object(s). Yet, they do not adapt well to changes in the user’s preferences. For instance, in fundus imaging, optic dis/cup weights more for glaucoma detection [29], while vessels are prioritized for assessing retinal vascular occlusion [26]. Even though these objects may appear within the same image, switching segmentation task from the optic disc/cup to vessels necessitates retraining the model. In real-world clinics where tasks are greatly diverse, creating separate models for individual tasks is practically challenging and computational demanding. Therefore, it is imperative to reinforce human-model interaction capacity by allowing users to flexibly adjust prompts for satisfying various clinical demands without the need for model retraining. For example, in intrapartum ultrasound as illuminated in Fig. 1(a), these properties would provide high flexibility for users to segment fetal head (FH) alone or FH&pubic symphysis (PS) at any time points $\{t_{1},t_{2}\}$ to achieve measurement of fetal head rotation or fetal position [7].
Confronted with these, we propose a novel Proxy Prompt (PP), which can be automatically generated from a “non-target” data (i.e., image/video frame of subjects other than the one under examination, such as from retrospective datasets) with a pre-annotated mask. This PP strategy is distinct from the existing prompting methods where prompting can only be made on “target” data, in a manual manner. As illustrated in Fig. 1(d), only one annotated image is required in using PP, tremendously streamlining workflow by waiving the prerequisite of providing separated prompts for every image/frame. Moreover, clinicians can freely switch segmentation tasks by adjusting the support-pair input anytime during examination without model retraining nor adopting different models, as shown in Fig. 1(b). Working in tandem with PP, we innovated a Proxy Prompt Generator (PPG) to reform SAM and SAM 2 for image and video data, respectively. Compared to SAM and SAM 2 in Fig. 2 (a&c), we employed high dimensional embedding from the PPG as prompts in Fig. 2 (b&d).
The core design of PPG lies in the novel Contextual Selective Module (CSM) and Contextual Colorization Module (CCM), which are dedicatedly configurated for automated prompting and high-level human-model interactions. First, CSM is introduced to enable adaptive selection of the most representative contextual information from “non-target” data for the “target” data, achieving cross-data guidance. Besides, CSM contains Vision Mamba, Bridge Unit, and Selective Map, implementing a 3-step selection process: (i) input-driven, (ii) object-guided, and (iii) target image/frame relevance selection to support both cross-video tasks and cross-image prompting. Second, CCM is devised to reinforce human-model interaction, enabling the model to interpret diverse user needs as indicated by different masks (e.g. single/multiple objects). This aim is achieved by leveraging dual-reverse cross-attention to enhance the representation of contextual embedding. Finally, the PP, effectively capturing specific object features, is generated. Such PPG-based strategy presents a simple yet efficient architecture to enhance clinical-friendliness of SAM and SAM 2, even in few-shot settings. Furthermore, with all original parameters frozen, our design can function as a flexible, plug-and-play module and can continuously adapt to the ever-evolving LVMs beyond SAM and SAM 2.
We conducted extensive experiments across several popular image and video datasets, validating the superior performance and stability of our proposed approach. Our main contributions are outlined below:
1. We propose a novel PP to enhance user-friendliness and precision of SAM and SAM 2 by equipping them with the capacity of automated prompting and high-level human-model interaction.
2. We devise CSM for adaptive selection of the most representative contextual information from “non-target” data to guide segmentation on “target” data, enabling effective cross-image/video prompting, waiving the need to execute prompting for every single target data, and minimizing experience-/expertise-derived variability in prompt quality.
3. We configurate CCM for enhancing the expressiveness of contextual embeddings to interpret and accommodate diverse clinical demands and preference of end users, thereby reinforcing model-human interactions.
4. Extensive experiments show that our model achieves state-of-the-art (SOTA) performance and is comparable to traditional segmentation models trained on full data volumes, even with only 16 image-mask pairs for SAM and SAM 2 training. Moreover, our strategy is of high potential to adapt to the iterative evolution of LVMs for medical tasks in the future.
2 Related Work
2.1 Adapting SAM to Medical Image Segmentation
Given the domain gap between natural and medical image datasets, various works [22, 37, 41, 27] have studied the application of SAM to medical image segmentation. Medical SAM (MedSAM) [22] fine-tunes the decoder of SAM using 1,570,263 medical image-mask pairs with bounding box prompts to adapt to medical tasks. Medical SAM Adapter (Med-SA) [37] efficiently fine-tunes SAM using adapter with point and box prompts. SAM Medical (SAMed) model [41] efficiently fine-tunes the image encoder of SAM using another technique—low-rank-based (LoRA) strategy [12]. AutoSAM [27] utilizes the gradients provided by a frozen SAM to train a new encoder, thus automatically extracts prompts from the image itself to achieve automatic segmentation. However, these SAM-based works on medical data can be roughly categorized into two types: one inherits the prompt design of SAM but without the automatic capability (MedSAM and Med-SA); the other can automatically segment objects but sacrifices the human-model interaction (SAMed and AutoSAM).
2.2 SAM without Manual Given Prompt
In addition to SAMed and AutoSAM, many other works also focus on enhancing the automatic prompting capability of SAM to improve user-friendliness. Self-prompting SAM [38] fine-tune a self-prompt unit to first provide coarse segmentations, from which they extract point/box prompts for SAM to obtain the final results. Personalization approach for SAM (PerSAM)  [42] identify the most similar point between the reference and test image as the prompt for SAM. Evidential prompt generation method (EviPrompt) [39] and ProtoSAM [1] both adapt the idea in PerSAM into medical image domain. EviPrompt fits the point prompts according to image similarity, while ProtoSAM ultilizes reference image/mask pair to obtain a coarse segmentation of target image, then extracts point or box as prompts required by SAM. However, these methods are still limited in simulating prompts such as points or boxes, which restricts their capability in the vessel-like branching structures due to the ambiguous instruction [22]. In contrast, our PP emphasizes the high-level embeddings as prompts, thereby guiding the model with deeper-level information to precisely segment such intricate objects.
2.3 Vision Backbone Based on Mamba
To extract contextual information for guiding model segmentation, the recently designed Mamba can serve as a potential choice for building vision backbone. Based on the state space model (SSM) [8], Mamba [9] boosts the development of SSM from the key aspects of “structure”. In terms of structure, it breaks the input-invariant feature of the conventional SSM layer and constructs an input-dependent SSM layer, enabling it to focus on the effective information in the input. Various studies [21, 15, 43, 23, 44] have migrated Mamba to the vision domain. Facial Expression Recognition-YOLO-Mamba (FER-YOLO-Mamba) [21] combines Mamba with attention to construct a dual-branch structure for facial expression detection. Vision Mamba-Denoising Diffusion Probabilistic Model (VM-DDPM) [15] introduces Mamba in the medical image synthesis domain, utilizing an SSM-CNN hybrid structure within the diffusion model. Vision Mamba (Vim) solely relies on the SSM to construct a vision backbone, selectively capturing key information in the input-dependent manner, making it highly suitable for handling high-resolution inputs [44], which is common in medical tasks.
Figure 3: Designed Proxy Prompt Generator for both SAM 2 and SAM. Our key designed focus on the Contextual Selective Module and the Contextual Colorization Module. The Encoder and Decoder refer to the original structures, which are frozen.
3 Method
Our proposed PPG, compatible with SAM and SAM 2, is illustrated in Fig. 3. Utilizing support image-mask pairs, this network not only enables auto-segmentation on target images or videos but also builds human-model interaction based on user-provided mask. For a target $\mathbf{X}$, we define a support set ${S}=\{{s}_{i}\}_{i=1}^{N}$ to assist in the segmentation of $N$ objects within the target $\mathbf{X}$. Each ${s}_{i}$ comprises $K$ image and mask pairs denoted as ${s}_{i}=\{(\mathbf{I}^{j},\mathbf{M}_{i}^{j})\}_{j=1}^{K}$, where each image $\mathbf{I}^{j}\in\mathbb{R}^{3\times H^{0}\times W^{0}}$ and the corresponding mask $\mathbf{M}_{i}^{j}\in\mathbb{R}^{H^{0}\times W^{0}}$. First, the target $\mathbf{X}$ and support images $\{\mathbf{I}^{j}\}_{j=1}^{K}$ are processed by a shared image encoder, with LoRA [12] applied to adapt with medical tasks, resulting in $\mathbf{F}_{x}\in\mathbb{R}^{C\times H\times W}$ and $\mathbf{F}_{\text{sup}}\in\mathbb{R}^{K\times C\times H\times W}$. In parallel, our Contextual Selective Module (CSM) processes the support pairs through a three-layer selection mechanism to identify the most valuable contextual information for target segmentation, thereby enabling cross-frame prompting (Sec. 3.1). Secondly, the contextual information $\mathbf{E}_{\text{ctx}}$ is refined within the Contextual Colorization Module (CCM), “coloring” target features with user specifications via dual-reverse cross-attention to achieve human-model interaction (Sec. 3.2). Finally, this refined high-level embedding—enriched with user-defined information and object-specific target feature—serves as a high-dimensional prompt for SAM or SAM 2 (Sec. 3.3). In summary, our method offers three main characteristics: cross-frame guidance, human-model interaction, and high-dimensional prompting via CSM and CCM. In the following sections, we detail each module step-by-step.
3.1 Contextual Selective Module
We leverage Vision Mamba [44]’s selective input capabilities, establish communication between different objects, and design a filtering mechanism to refine the contextual embedding. As illustrated in Fig. 3, support pairs $\{{s}_{i}\}_{i=1}^{N}$ are initially processed by the Vision Mamba encoder to extract feature $\mathbf{V}$ with input-adaptive parameters, marking the first selective step. Secondly, within the designed Bridge Unit, we enable inter-object communication across the concatenated features, thereby selectively extracting object-level feature $\mathbf{A}_{\text{agg}}$. Thirdly, features ($\mathbf{F}_{\text{x}}$) and ($\mathbf{F}_{\text{sup}}$) are used to compute a $\mathbf{Selective}$ map. Each row in this map representing a given patch from a support image will positively influence a specific patch in $\mathbf{X}$. This selective map subsequently filters the aggregated information $\mathbf{A}_{\text{agg}}$ to derive the most valuable information as the contextual embedding $\mathbf{E}_{\text{ctx}}$. The following sections will detail each component of this pipeline.
First Selection Step: Vision Mamba. We use the encoder of Vision Mamba (Vim)  [44] as the contextual encoder, as its input-dependent selection mechanism is well-suited for learning similar modality features, thereby enhancing prompt effectiveness. This mechanism enables parameters (i.e., $\mathbf{\overline{A}}$, $\mathbf{\overline{B}}$ and $\mathbf{C}$) influencing interactions within the image to adapt as functions of the input, allowing features to be extracted based on the input itself. Specifically, Vim first flattens the input 2-D image into patches and form the token sequence $x$, which is progressively transformed into the output token sequence $y$ according to the following formula, ultimately being encoded by us into the feature matrix.
$h_{t}=\mathbf{\overline{A}}h_{t-1}+\mathbf{\overline{B}}x_{t}$ (1)
$y_{t}=\mathbf{C}h_{t}$ (2)
where $t$ represents the timestep and $h$ denotes the latent state. The parameters $\mathbf{\overline{A}}$, $\mathbf{\overline{B}}$ and $\mathbf{C}$ are generated depending on the input $x$, rather than being input-invariant, thus allowing us to selectively encode the input image as our first step in the selection strategy. For more details, refer to [44].
Given the support set ${S}=\{{s}_{i}\}_{i=1}^{N}$, we fix a value $K$ such that each ${s}_{i}=\{(\mathbf{I}^{j},\mathbf{M}_{i}^{j})\}_{j=1}^{K}$ provides $K$ support pairs $(\mathbf{I}^{j},\mathbf{M}_{i}^{j})$. For the $j$-th pair $\{(\mathbf{I}^{j},\mathbf{M}_{i}^{j})\}_{i=1}^{N}$ among the $N$ support subset ${s}_{i}$, each masks $\mathbf{M}_{i}^{j}$ correspond to one same image $\mathbf{I}^{j}$. As a result, $\mathbf{I}^{j}$ is concatenated with $\mathbf{M}_{i}^{j}$ and fed into the contextual encoder, resulting in $N$ feature matrices $\mathbf{V}\in\mathbb{R}^{K\times C^{v}\times H\times W}$, where $C^{v}$ indicates the channel of $\mathbf{V}$. The process of extracting the feature matrix $\mathbf{V}$ can be breifly represented by the following equation:
[…]
Second Selection Step: Bridge Unit. Subsequently, in the Bridge Unit, we initially duplicate $N$ times the obtained feature $\mathbf{F}_{\text{sup}}$ and concat with $\mathbf{V}$ of each object along the channel dimension to further aggregate the features. Most importantly, we employ convolutional block attention module(CBAM) [35] along the concatenated channel dimension to facilitate implicit inter-target communication, allowing the model to select key features across multiple objects, thereby enhancing cross-target contextual understanding. Additionally, two ResBlocks  [10] are used to prevent further feature dimension expansion. After the final block, the feature is flattened to form $\mathbf{A}_{agg}$. This process is summarized as follows:
$\mathbf{F}_{\text{cat}}=\text{concat}(\mathbf{F}_{\text{sup}},\mathbf{V})$ (4)
$\mathbf{A}_{agg}=\text{ResBlock}_{2}(\text{CBAM}(\text{ResBlock}_{1}(\mathbf{F}_{\text{cat}})))$ (5)
where $\mathbf{F}_{\text{cat}}\in\mathbb{R}^{K\times N\times(C^{v}+C)\times H\times W}$ and $\mathbf{A}_{agg}\in\mathbb{R}^{N\times C\times(K\times H\times W)}$ represents the aggregated features of the support set.
Third Selection Step: Selective Map. In order to select the key information for the target features, we compute the selective map based on $\mathbf{F}_{x}$ and $\mathbf{F}_{\text{sup}}$. First, $\mathbf{F}_{x}$ and $\mathbf{F}_{\text{sup}}$ are flattened into $\mathbf{\hat{F}}_{x}\in\mathbb{R}^{C\times(H\times W)}$ and $\mathbf{\hat{F}}_{\text{sup}}\in\mathbb{R}^{C\times(K\times H\times W)}$, respectively. Then, the map is calculated through matrix multiplication, following  [5] and using the following equation:
$\mathbf{Selective}=\frac{2\cdot(\mathbf{\hat{F}}_{\text{sup}}^{T}\cdot\mathbf{\hat{F}}_{x})-\mathbf{\hat{F}}_{\text{sup}}^{2}}{\sqrt{C}}$ (6)
where $C$ is the channel dimension of $\mathbf{\hat{F}}_{\text{sup}}$ and $\mathbf{Selective}\in\mathbb{R}^{(K\times H\times W)\times(H\times W)}$. Afterward, the computed $\mathbf{Selective}$ is normalized using the softmax to ensure that the contribution values conform to a probability distribution. The resultant information matrix can be represented as follows:
$\mathbf{E}_{\text{ctx}}=\mathbf{A}_{agg}\cdot Softmax(\mathbf{Selective})$ (7)
where $\mathbf{E}_{\text{ctx}}$ serves as the contextual (ctx) embedding of the key information from the support set, and $\mathbf{E}_{\text{ctx}}\in\mathbb{R}^{N\times C\times(H\times W)}$.
3.2 Contextual Colorization Module
The CCM is proposed to interpret user intent from the contextual embedding derived from the support set, enabling human-model interaction. Conceptually, this process is similar to “colorizing” the target image based on the support mask. Unlike the CSM, which focuses on selecting contextually representative information tailored to the target image, this module emphasizes dynamically refining target features on the ground of the contextual embedding via cross-attention, thereby facilitating a deeper understanding of the user’s segmentation intent. Before entering the module, $\mathbf{\hat{F}}_{\text{x}}$ is duplicated $N$ times to align its dimensions with $\mathbf{E}_{\text{ctx}}$.
The CCM consists of four identical blocks, one of which is detailed in Fig. 3. In each block, $\mathbf{\hat{F}}_{\text{x}}$ and $\mathbf{E}_{\text{ctx}}$ are each passed through a learnable projection layer to reduce their dimensions to $\mathbf{\hat{F}}_{\text{x}}\in\mathbb{R}^{N\times C\times((H\times W)/2)}$ and $\mathbf{E}_{\text{ctx}}\in\mathbb{R}^{N\times C\times((H\times W)/2)}$, respectively. Subsequently, $\mathbf{E}_{\text{ctx}}$ is integrated into the target image features to guide the model in identifying specific regions expected to be “colored”. This integration occurs by adding cross-attention-processed information back into $\mathbf{\hat{F}}_{\text{x}}$, followed by a feed-forward layer (FFN) [30] for contextual reasoning, yielding $\mathbf{\hat{F}}_{\text{x\_next}}$, which then serves as the next target feature input for the next block. This process is represented as follows:
$\mathbf{\hat{F}^{{}^{\prime}}}_{\text{x}}=\text{Add\&Norm}(\text{Cross Attention}(\mathbf{\hat{F}}_{\text{x}},\mathbf{E}_{\text{ctx}}))$ (8)
$\mathbf{\hat{F}}_{\text{x\_next}}=\text{Add\&Norm}(\text{FFN}(\mathbf{\hat{F}^{{}^{\prime}}}_{\text{x}}))$ (9)
Subsequently, $\mathbf{E}_{\text{ctx}}$ reads from the updated target image features $\mathbf{\hat{F}}_{\text{x\_next}}$ through a reversed cross-attention layer to pinpoint features essential for matching object information in the contextual embedding. In the followed FFN, the context embedding, equipped with object-specific feature representation, undergoes further enhancement to strengthen its segmentation guidance capability. The resulting $\mathbf{E}_{\text{ctx\_next}}$ then serves as the next context embedding for the next block. This process can be expressed as follows:
$\mathbf{E}^{{}^{\prime}}_{\text{ctx}}=\text{Add\&Norm}(\text{Cross Attention}(\mathbf{E}_{\text{ctx}},\mathbf{\hat{F}}_{\text{x\_next}}))$ (10)
$\mathbf{E}_{\text{ctx\_next}}=\text{Add\&Norm}(\text{FFN}(\mathbf{E}^{{}^{\prime}}_{\text{ctx}}))$ (11)
After passing through the final block, the iteratively refined context embedding fully comprehends the user-defined segmentation intent from the support set, as well as the corresponding image features derived from the target image. This enhanced context embedding subsequently serves as $N$ prompts $\mathbf{P}$ for the SAM or SAM 2 model, where $\mathbf{P}\in\mathbb{R}^{N\times((H\times W)/16)\times C}$.
3.3 Loss Function
For SAM, the image features $\mathbf{F}_{x}$ extracted from the target image $\mathbf{X}$ is fed into the Decoder with the prompt $\mathbf{P}$ to generate the final output, represented by the following equation:
$\text{Output}=\text{Decoder}(\mathbf{F}_{x},\mathbf{P})$ (12)
where $\text{Output}\in\mathbb{R}^{N\times H^{0}\times W^{0}}$. For SAM 2, our approach directly inputs the high-dimensional vector $\mathbf{P}$ into memory attention as prompt embeddings, thus $\mathbf{P}$ also can be considered as memory extracted across video frames. For loss settings, we simply follow the design of SAM  [16] to supervise the mask prediction with a Dice loss  [24], using the following formula:
$\mathcal{L}=1-\frac{2\sum_{i=1}^{D}p_{i}g_{i}}{\sum_{i=1}^{D}p_{i}+\sum_{i=1}^{D}g_{i}}$ (13)
where $p_{i}$ and $g_{i}$ represent the probability of pixel $i$ in the predicted mask and the label of pixel $i$ in the ground truth mask, respectively. The number of pixels involved in the computation of the Dice loss is denoted by $D$.
Method Training Images Color Fundus Photography Ultrasound Avg. Dice $\uparrow$ (%)
REFUGE2-Disc REFUGE2-Cup STARE-Vessel FPA-PS FPA-FH
$Upper^{*}$ $2~{}\text{x}~{}10^{3}$ 93.7 83.5 65.5 82.1 91.7 83.3
[…]
Med-SA (1-point) 16 85.6 83.0 45.4 71.3 81.4 73.3
Med-SA (box) 16 86.8 82.6 37.1 71.8 83.1 72.3
SAMed 16 86.9 84.3 15.8 66.3 87.0 68.1
AutoSAM 16 87.9 84.6 73.8 78.0 79.1 80.7
Ours 16 88.0 85.5 81.6 85.3 89.3 85.9
Table 1: Comparison of our model with SAM-based SOTAs in Dice Score (%) on image datasets, including train-free models (denoted as ${}^{\text{TF}}$), efficient fine-tuned models, and models trained on full data (denoted as $Upper^{*}$). ‘Training Images’ denotes the average number of images used for training to segment five objects, with dash(–) denoting inapplicability. Gray indicates the data were used in model pre-training. The $\bf best$ and $\underline{\text{second-best}}$ results (excluding $Upper$) are bolded and underlined, respectively, showcasing our SOTA performance among various models under same few-shot settings.
Method Training data Avg. Dice $\uparrow$ (%)
With Label W/O. Label
$Upper^{*}$ 1134 (100%) 0 (0%) 88.1
DTC 58 (5.1%) 1076 (94.9%) 63.4
MLB-Seg 58 (5.1%) 1076 (94.9%) 78.3
nnUnet 58 (5.1%) 0 (0%) 84.1
Med-SA 58 (5.1%) 0 (0%) 76.0
SAMed 58 (5.1%) 0 (0%) 85.9
AutoSAM 58 (5.1%) 0 (0%) 70.1
Ours 58 (5.1%) 0 (0%) 87.4
Table 2: Comparison of our model against various SOTAs on MRI dataset, including semi-supervised models [20, 34], SAM-based models, traditional segmentation models, and 3D DSD-FCN [31] trained on the full dataset (denoted as $Upper^{*}$). ‘Training Data’ indicates the number of slices used for training and their percentage of the total dataset.
4 Experiment
4.1 Datasets and Implementation Details
We conducted extensive experiments on five widely used publicly available datasets, including REFUGE2 [6], STARE [11], FH-PS-AOP(FPA) [14], PROMISE12 [18] (3D MRI) and JNU-IFM [19] (Video). Among these, STARE [11] and PROMISE12 [18] contain only one single object, while the other datasets contain multiple segmentation objects. Further introduction in the supplementary.
Method Prompt Preload
Capable Avg.$\uparrow$
(%) Std.$\downarrow$
[…]
$Upper$ Mask 89.6 - -
SAM2 Box 78.6 - -
SAM2 Point 25.1 - -
MedSAM2 Mask 63.3 10.9 29.1
Ours Mask 80.9 0.3 1.0
Table 3: Comparison of our model with SAM2-based SOTAs on the video dataset, reporting Average Dice, standard deviation Dice, and max-min Dice difference. $Preload~{}Capable$ indicates whether the prompt can be pre-prepared for real-time use. Background shading from light to dark green represents increasing real-time usability, while SAM2-Mask, being nearly impractical, is treated as $Upper$ and marked in Gray. The $\bf best$ and $\underline{\text{second-best}}$ results (excluding $Upper$) are bolded and underlined, respectively.
Due to the limited training data, we freeze the pre-trained parameters of SAM to avoid overfitting. We applied LoRA [12] to the original encoder and decoder for efficient fine-tuning on medical datasets. Model optimization was performed using the Stochastic Gradient Descent (SGD) optimizer, with a momentum of 0.9, a learning rate of 0.01, and a weight decay of 0.0005. For image and video datasets, we randomly selected 16 patient-level images or videos for training, with the remaining samples used for testing. For 3D data, 58 labeled slices from 3 patients were used for training, while data from 10 additional patients were reserved for testing. Further details in the supplementary.
Figure 4: Visualization comparison results of nine models across five objects.
4.2 Comparison with SOTA on Image Dataset
We extensively evaluated diverse comparison methods, including train-free foundation models, models efficiently fine-tuned on the specific-dataset, and traditional models ($Upper$). Details are in the supplementary. MedSAM, pretrained on a large medical dataset, was evaluated on both its seen FPA dataset and the unseen REFUGE2 dataset. For trainable models (Med-SA, SAMed, AutoSAM, and ours), we trained each on the same 16 images until convergence.
As shown in Tab. 1, MedSAM and AutoSAM showed strong performance in both train-free and few-shot settings, probably due to pre-training on large medical datasets or object-specific training. However, MedSAM’s performance dropped significantly on unseen datasets and struggled with intricate structure such as vessel in the STARE dataset. AutoSAM, on the other hand, underperformed on ultrasound datasets, possibly due to the challenges such as ultrasound imaging-associated speckle noise. Our method achieved SOTA performance under limited conditions, even comparable with traditional SOTAs trained on the full dataset. Fig. 4 shows that while SAMed and AutoSAM performed relatively well on REFUGE2 and FPA, both still struggled with boundary omissions. On STARE, MedSAM’s box prompt led to segmentation ambiguity. Across all datasets, our method delivered consistently superior and stable results.
4.3 Comparison with SOTA on MRI Dataset
Although not specifically designed for 3D data, our model shows strong potential. We evaluated it on the 3D MRI PROMISE12 dataset [18] (details in the supplementary). As shown in Tab. 2, it achieved the highest few-shot performance (87.4%), surpassing all SAM-based models and the traditional nnUnet under the same conditions. Notably, despite not leveraging any additional unlabeled data, our method surpassed the best semi-supervised model (MLB-Seg [34], 78.3%). Furthermore, we achieved results comparable to the fully-supervised $Upper^{*}$ (88.1%), with only a 0.7% gap, showcasing its promise for 3D segmentation.
4.4 Comparison with SOTA on Video Dataset
We conducted a comprehensive comparison on the JNU-IFM dataset across four models: MedSAM2, SAM2-box, SAM2-point, and Ours. Since Ours and MedSAM2 allow pre-prepared prompts from non-target data, they offer the simplest real-time usability, followed by SAM2-point (one-click), and lastly SAM2-box. Although mask prompts provide the most detailed information, they are impractical for a single operator in real-time and thus serve as our $Upper$. To assess the stability of MedSAM2 and Ours, we conducted five trials on the same test set, each using a different support pair randomly drawn from five pre-excluded videos. Further experimental details are in supplementary. As shown in Sec. 4.1, our method achieves the highest average Dice, improving by 17.6% over MedSAM2, which uses a simple prompt strategy, and by 2.3% over SAM2-box, which requires the most demanding prompt strategy. Further, our method yields comparable performance to the $Upper$, and presents excellent stability indicated by 0.3% Std. Dice (far greater than MedSAM2 of 10.9% Std. Dice and 29.1% fluctuation in Dice).
CSM CCM Avg. Dice $\uparrow$ (%) std. Dice $\downarrow$ (%)
85.9 2.6
85.2 3.1
84.3 3.0
84.0 3.1
Table 4: Ablation experiments on different PPG modules.
Figure 5: Selective Map visualization on support images with different mean square errors (MSE) to the target image.
4.5 Ablation Studies
We analyze the impact on PP-SAM from multiple aspects, including model architecture, pretrained weights, prompt quality and quantity during inference, and training data size. Experiment settings and details are in the supplementary.
Effectiveness of each model component. The results in Tab. 4 highlight the contributions of both CSM and CCM to the optimal model performance. When CCM was removed, the segmentation performance std increased significantly (2.6% → 3.1%), indicating inconsistencies in multi-object segmentation. Removing CSM led to an overall degradation in segmentation quality, with the average Dice score decreasing by 1.6%. When CSM and CCM were simultaneously omitted, the model achieved the lowest Avg. Dice (84.0%) with the highest std. Dice (3.1%), confirming that CSM and CCM collectively enhance both segmentation accuracy and stability.
Visualization of selective map effectiveness across different support images. Fig. 5 shows that the Selective Map directs the target image to focus on relevant anatomical structures in the support image while suppressing background interference. Across varying support image similarity, from S1 (MSE = 0) to S2 (MSE = 3010.1), it consistently highlights semantically relevant regions, ensuring stable segmentation (Avg. Dice: 95.42%–95.53%).
Effect of encoder architectures in CSM shown in Tab. 11. Vision Transformer achieves better results than ResNet-50 (Avg. Dice: 84.5% vs. 83.9%), which may be attributed to significantly larger number of trainable parameters (89.0M vs. 8.5M). Notably, Vision Mamba attains the highest performance (Dice: 85.9%; IoU: 76.3%) despite having the lowest parameter count (7.9M).
Encoder in CSM Turnable Param (M) Avg. Dice $\uparrow$ (%) Avg. IoU $\uparrow$ (%)
Vision Mamba 7.9 85.9 76.3
Vision Transformer 89.0 84.5 74.0
ResNet-50 8.5 83.9 72.6
Table 5: Ablation on different encoder architectures in CSM.
Ablation on pretrained foundation models. Tab. 6 shows that our Proxy Prompt consistently enhances SAM-based models, including MedSAM, which was pretrained on extensive medical data. Surprisingly, MedSAM with our prompt underperforms compared to SAM with the same prompt, likely due to its strong reliance on Box prompts during pretraining, limiting the adaptability to our prompt. Deep analysis is in the supplementary.
Method Pretrained Weight Prompt Method Avg. Dice $\uparrow$ (%)
MedSAM MedSAM-ViT-B Box 74.5
Proxy Prompt 81.3 (+6.8)
SAM SAM-ViT-B Box 56.1
Proxy Prompt 85.1 (+29.0)
SAM-ViT-H Box 60.3
Proxy Prompt 85.9 (+25.6)
Table 6: Performance with different pretrained foundation models.
Influence of prompt quality via support mask variations. We evaluate six levels of support mask quality, recording each mask’s Dice score as Prompt Dice in Fig. 9. The results shows segmentation improves (visualized by intensifying red shades) with prompt quality, dropping sharply only in $Empty$ level. Our model remains robust regardless of whether masks come from normal users or senior doctors. Line chart is in the supplementary.
Empty Low Medium High SD NU
Prompt Dice (%) 00.0 66.6 81.6 100 88.3 74.8
Output Dice (%) 47.9 77.8 86.5 86.8 86.6 86.5
Table 7: Model performance across prompt qualities, including senior doctor (SD) and normal user (NU). Prompt Dice reflects support mask quality, while Output Dice measures performance.
Figure 6: The boxes show performance for each $M$ with varying $K$, while in the rightmost plots ($G<M-1$), higher peaks refer to better performance and narrower ranges reflect improved stability. Optimum is achieved when both $M$ and $K$ are maximized.
Impact of training data size $M$ and support pair quantity $K$ in inference. Fig. 11 shows that performance improves with $M$. For each $M$, Dice increases with $K$ until $K$ exceeds $M-1$, after which a slight decline may occur. The $M=16$ ridge plot indicates that as $K$ increases, prediction variance decreases significantly. Overall, both training dataset size $M$ and support size $K$ influence model performance, which stabilizes at its peak when both are maximized. Experiment details are in the supplementary.
5 Conclusion
We present the PPG, a pluggable framework designed to endow SAM and SAM 2 with auto-prompting and enhanced their interactive capabilities. Our PPG employs a CSM to extract contextual information from non-target data and a CCM to interpret user intentions, enabling it to generate efficient PP that effectively guides the model output to meet user needs.